# MemBench: Memorized Image Trigger Prompt Dataset for Diffusion Models

### v1

@@ 3.4 Dataset Construction Leveraing MCMC
[…]
@@
This stage aims to augment memorized image trigger prompts. We leverage the prompts found in the previous stage or those provided by Webster [26] as the prior, $\pi_{0}$. In this process, we employ a slightly modified algorithm to enhance diversity. Instead of running a single chain for one prompt, we run $n$ separate chains for each word position in an $n$-length sentence, treating each position as the first updating index in Gibbs sampling. This method ensures a varied exploration of the prompt space. We then save the top 100 prompts with the highest $D_{\bm{\theta}}(\cdot)$. For these prompts, we generate 10 images each and filter the prompts that generate images with the highest SSCD to the dataset. The detailed process is provided in Supplementary Material.
@@ 4.1 Dataset Analysis
Stable Diffusion 1 Stable Diffusion 2
Trigger Prompt # Mem. Image # Trigger Prompt # Mem. Image #
Webster [26] 325 114 210 25
MemBench (Ours) 3000 154 1500 55
Table 1: Comparison of the number of memorized images and trigger prompts provided in each dataset. Our dataset is near 9 times larger in terms of the number of trigger prompts.
@@
Table 1 presents the number of memorized images and trigger prompts obtained using the methodology described in Section 3. For both Stable Diffusion 1 and 2, the number of prompts has increased more than fivefold for each model and more than 9 times in total compared to those reported by Webster [26]. Additionally, the number of memorized images included in the dataset has also increased, with Stable Diffusion 2 showing an increase of over twofold. The composition of the images included in MemBench is shown in Figure 1, illustrating that the memorized images encompass a substantial number of commercial product images and human images. It also includes artwork such as brand logos.
@@ 4.2 Efficacy of Memorized Image Trigger Prompt Searching
In this section, we validate the efficiency of our method in discovering memorized images without access to $\mathcal{D}_{train}$. The task of finding memorized image trigger prompts without $\mathcal{D}_{train}$ is defined as follows: without any prior information, the method must automatically find trigger prompts that induce memorized images. This involves: 1) selecting candidate prompts, 2) generating 100 images for each candidate prompt, 3) applying DBSCAN [3] clustering with SSCD to get candidate images and using a Reverse Image Search API to verify those images’ presence on the web. To the best of our knowledge, this task is novel, so we provide naive baselines. As the first baseline, we perform a greedy search by measuring $D_{\theta}$ for all prompts in the prompt dataset and selecting the top 200 prompts as candidate prompts. For the prompt dataset, we leveraged DiffusionDB, which contains 13M prompts collected from diffusion model users. Additionally, we provide three other baselines, all of which are algorithms that solve optimization problems in the prompt space. For two of these baselines, we adapt ZeroCap [25] and ConZIC [30], methods designed to maximize the CLIP Score for zero-shot image captioning, by replacing their objective function with $D_{\theta}$. Similarly, we also adapt PEZ [28] by substituting its objective function with $D_{\theta}$ to serve as another baseline. For each of these three methods, we conducted 200 iterations and obtained prompts. For our method, we performed 200 MCMC runs with 150 iterations each, and selected the resulting prompts as candidate prompts. For more detailed implementation, please refer to the Supplementary Material.
The results are shown in Table 2. Our method significantly outperforms other methods. Comparing with baselines [1, 26] that levrerage LAION itself is challenging, as the dataset is no longer available and the elements for reimplementation are omitted in the corresponding papers. However, as mentioned in Section 2, their memory-inefficient and computationally intensive methods provided only few memorized images and trigger prompts. Please refer to the Supplementary Material for the performance of augmenting the memorized image trigger prompts through MCMC.
@@ 5 MemBench: Metrics, Scenarios and Reference Performance
We present rigorous metrics for correctly evaluating mitigation methods, which include similarity score, Text-Image alignment score, and quality score. Firstly, following previous works, we adopt SSCD [14] as the similarity score and measure max SSCD between a generated image using trigger prompt and memorized images. In detail, if a prompt ${\mathbf{p}}^{*}$ triggers images $\{{\mathbf{x}}_{1}^{*},...,{\mathbf{x}}_{k}^{*}\}$ included in $\mathcal{D}_{train}$, we measure $\max_{{\mathbf{x}}\in\{{\mathbf{x}}_{1}^{*},...,{\mathbf{x}}_{k}^{*}\}}SSCD({\mathbf{x}}({\mathbf{p}}^{*},{\bm{\epsilon}}_{\theta}),{\mathbf{x}})$. Secondly, we adopt CLIP Score [7] to measure Text-Image alignment between prompt and generated images. Lastly, We adopt an Aesthetic Score [19] as the image quality score. While previous works did not measure image quality scores, we observed issues shown in Figure 2. When memorization mitigation methods are applied, we observed that image quality degrades, the rich context generated by the diffusion model is destroyed, or distorted images are formed. An ideal memorization mitigation method should be able to preserve the image generation capabilities of the diffusion model.
@@
To ensure that memorization mitigation methods can be generally applied to diffusion models, we provide two scenarios: the memorized image trigger prompt scenario and the general prompt scenario. First, the memorized image trigger prompt scenario evaluates whether mitigation methods can effectively prevent the generation of memorized images. This scenario uses the memorized image trigger prompts we identified in Section 3. In this scenario, we generate 10 images for each trigger prompt and measure the Top-1 SSCD and the mean values of the Top-3 SSCD. We also measure the proportion of images with SSCD exceeding 0.5. For CLIP Score and Aesthetic Score, we calculate the average value across all generated images. Second, the general prompt scenario ensures that the performance of the diffusion model does not degrade when using prompts other than trigger prompts. We leverage the COCO [11] validation set as general prompts. In this scenario, images are generated once per prompt, and the average CLIP Score and Aesthetic Score are measured.
@@
[…]
@@ 6 Deeper Analysis into Image Memorization
In this section, we provide a deeper analysis of the memorized images and trigger prompts in MemBench. Firstly, we have found that Stable Diffusion regenerates commercial products currently on sale. While the possibility that diffusion models could memorize commercial images has been suggested [1, 20], we are the first to confirm this. Unlike the previous studies [1, 26] that used image retrieval from LAION to find memorized images, we leverage a Reverse Image Search API to find those, which enable us this verification. As shown in Figure 3.b, Stable Diffusion replicates images of commercially available shirts when given a specific prompt. Figure 3.d further illustrates the replication of layouts; for a commercially sold carpet, all layouts have been reproduced.
[…]
@@ 7 Evaluation of Image Memorization Mitigation Method on MemBench
Base 0.737 0.712 0.638 0.299 5.28 0.321 5.37
Reference Performance (API search) 0.210 - - 0.329 - - -
RNA[21] n = 1 0.629 0.587 0.435 0.290 5.22 0.314 5.34
n = 2 0.550 0.500 0.318 0.286 5.19 0.310 5.33
n = 3 0.473 0.426 0.240 0.281 5.16 0.307 5.30
n = 4 0.427 0.379 0.193 0.278 5.14 0.304 5.29
n = 5 0.383 0.337 0.148 0.273 5.11 0.302 5.28
RTA [21] n = 1 0.639 0.597 0.449 0.289 5.23 0.316 5.34
n = 2 0.540 0.496 0.319 0.284 5.22 0.314 5.32
n = 3 0.464 0.418 0.232 0.281 5.19 0.310 5.29
n = 4 0.426 0.370 0.183 0.277 5.19 0.306 5.27
n = 5 0.368 0.326 0.136 0.273 5.16 0.302 5.26
Wen et al. [29] l = 12 0.625 0.580 0.415 0.292 5.23 0.321 5.37
l = 9 0.563 0.506 0.292 0.286 5.19 0.321 5.37
l = 6 0.466 0.395 0.145 0.276 5.16 0.321 5.37
l = 5 0.420 0.347 0.100 0.274 5.15 0.321 5.37
l = 4 0.354 0.287 0.059 0.269 5.13 0.321 5.37
Ren et al. [16] c = 1.0 0.388 0.336 0.120 0.275 5.21 0.316 5.33
c = 1.1 0.379 0.328 0.103 0.271 5.20 0.313 5.31
c = 1.2 0.373 0.322 0.088 0.267 5.19 0.309 5.28
c = 1.3 0.370 0.318 0.080 0.263 5.18 0.304 5.26
Table 3: Performance evaluation of image memorization mitigation methods in MemBench. Please refer to Supplementary Material for the details of hyper-paramters.
In this section, we evaluate image memorization mitigation methods on our MemBench in Stable Diffusion 1. For results of Stable Diffusion 2, please refer to the Supplementary Material.
@@
We use Stable Diffusion 1.4 as the base model. The image memorization mitigation methods evaluated include: 1) RTA [21], which applies random token insertion to the prompt, 2) RNA [21], which inserts a random number between $[0,10^{6}]$ into the prompt, 3) method proposed by Wen et al. [29] that applies adversarial attacks to text embeddings, and 4) method proposed by Ren et al. [16] that rescales cross-attention. Image generation is performed using the DDIM [22] Scheduler with a guidance scale of 7.5 and 50 inference steps.
@@ Results.
We present the experimental results in Table 3. As shown in Table 3, for all methods, lowering the SSCD significantly reduces both the CLIP Score and Aesthetic Score. This indicates a degradation in Text-Image alignment and image quality. Specifically, to reduce the proportion of generated images with an SSCD exceeding 0.5 below 20%, all methods need to lower the CLIP Score by more than 0.02. Given the scale of the CLIP Score, this level of reduction suggests that the generated images are barely related to the given prompt.
When comparing methods, RNA [21] and RTA [21] are the most effective in lowering SSCD while preserving the CLIP Score and Aesthetic Score for the trigger prompts. However, these methods are impractical as they also significantly reduce the CLIP Score and Aesthetic Score in the general prompt scenario, COCO. Conversely, the method proposed by Wen et al. [29] maintains the CLIP Score and Aesthetic Score in the general prompt scenario, indicating its suitability for practical applications. The method proposed by Ren et al. [16] shows a substantial reduction in both CLIP Score and Aesthetic Score, even with minimal hyperparameter settings. Furthermore, images obtained through API queries with trigger prompts can achieve an SSCD as low as 0.210 while maintaining high CLIP Score. This indicates that the performance of current mitigation methods is still lacking.

### v2

@@ 3.4 Dataset Construction Leveraing MCMC
This stage aims to discover new memorized images. The sentence is initialized with sentence of length $n$ [MASK] token, i.e. ${\mathbf{p}}_{0}=$[MASK], [MASK], … , [MASK]}. We then employ Algorithm 1 initialized with ${\mathbf{p}}_{0}$ to obtain the candidate prompt. Similar to the conventional approach (Carlini et al., 2023) to extract train images, we then generate 100 images for this prompt and leverage DBSCAN (Ester et al., 1996) clustering algorithm with SSCD (Pizzi et al., 2022) to extract images forming at least 20 nodes. Those images are employed to Reverse Image Search API to find train image sources and human verification is conducted.
@@
This stage aims to augment memorized image trigger prompts. We leverage the prompts found in the previous stage or those provided by Webster (2023) as the prior, $\pi_{0}$. In this process, we employ a slightly modified algorithm to enhance diversity. Instead of running a single chain for one prompt, we run $n$ separate chains for each word position in an $n$-length sentence, treating each position as the first updating index in Gibbs sampling. This method ensures a varied exploration of the prompt space. We then save the top 100 prompts with the highest $D_{\bm{\theta}}(\cdot)$. We retained all prompts generated during the MCMC sampling process and then selected 20 augmented prompts per original prompt, considering diversity. The detailed process is provided in Appendix.
@@ 4.1 Dataset Analysis
Stable Diffusion 1 Stable Diffusion 2 DeepFloydIF Realistic Vision
Trigger Prompt # Mem. Image # Trigger Prompt # Mem. Image # Trigger Prompt # Mem. Image # Trigger Prompt # Mem. Image #
Webster (2023) 325 111 210 25 162 17 354 119
MemBench 3000 151 1500 55 309 51 1352 148
Table 1: Comparison of the number of memorized images and trigger prompts provided in each dataset. Our dataset is significantly larger in terms of the number of trigger prompts across all models. Please note that images sharing the same layout, as depicted in Figure 4, have been counted as a single image.
@@
Table 1 presents the number of memorized images and trigger prompts obtained using the methodology described in Section 3. For both Stable Diffusion 1 and 2, the number of prompts has increased more than fivefold for each model and more than 9 times in total compared to those reported by Webster (2023). The number of memorized images included in the dataset has also increased, with Stable Diffusion 2 showing an increase of over twofold. Additionally, we provide memorized images and trigger prompts for DeepFloydIF (Shonenkov et al., 2023), which has a cascaded structure, and Realistic Vision (CivitAI, 2023), an open-source diffusion model. For these two models as well, we provide a greater number of memorized images and trigger prompts than Webster et al. We have also applied our algorithm to the more recent model, Stable Diffusion 3 (Esser et al., 2024). Please refer to Appendix for the results. The composition of the images included in MemBench is shown in Figure 1, illustrating that the memorized images encompass a substantial number of commercial product images and human images. It also includes artwork such as brand logos.
@@ 4.2 Efficacy of Memorized Image Trigger Prompt Searching
In this section, we validate the efficiency of our method in discovering memorized images without access to $\mathcal{D}_{train}$. The task of finding memorized image trigger prompts without $\mathcal{D}_{train}$ is defined as follows: without any prior information, the method must automatically find trigger prompts that induce memorized images. This involves: 1) selecting candidate prompts, 2) generating 100 images for each candidate prompt, 3) applying DBSCAN (Ester et al., 1996) clustering with SSCD to get candidate images and using a Reverse Image Search API to verify those images’ presence on the web. To the best of our knowledge, this task is novel, so we provide naive baselines. As the first baseline, we perform a greedy search by measuring $D_{\theta}$ for all prompts in the prompt dataset and selecting the top 200 prompts as candidate prompts. For the prompt dataset, we leveraged DiffusionDB, which contains 13M prompts collected from diffusion model users. Additionally, we provide three other baselines, all of which are algorithms that solve optimization problems in the prompt space. For two of these baselines, we adapt ZeroCap (Tewel et al., 2022) and ConZIC (Zeng et al., 2023), methods designed to maximize the CLIP Score for zero-shot image captioning, by replacing their objective function with $D_{\theta}$. Similarly, we also adapt PEZ (Wen et al., 2024a) by substituting its objective function with $D_{\theta}$ to serve as another baseline. For each of these three methods, we conducted 200 iterations and obtained prompts. For our method, we performed 200 MCMC runs with 150 iterations each, and selected the resulting prompts as candidate prompts. For more detailed implementation, please refer to Appendix.
The results are shown in Table 2. Our method significantly outperforms other methods. Comparing with baselines (Carlini et al., 2023; Webster, 2023) that leverage LAION itself is challenging, as the dataset is no longer available and the elements for reimplementation are omitted in the corresponding papers. However, as mentioned in Section 2, their memory-inefficient and computationally intensive methods provided only few memorized images and trigger prompts. Please refer to Appendix for the performance of augmenting the memorized image trigger prompts through MCMC.
@@ 5 MemBench: Metrics, Scenarios and Reference Performance
We present rigorous metrics for correctly evaluating mitigation methods, which include similarity score, Text-Image alignment score, and quality score. Firstly, following previous works, we adopt SSCD (Pizzi et al., 2022) as the similarity score and measure max SSCD between a generated image using trigger prompt and memorized images. In detail, if a prompt ${\mathbf{p}}^{*}$ triggers images $\{{\mathbf{x}}_{1}^{*},...,{\mathbf{x}}_{k}^{*}\}$ included in $\mathcal{D}_{train}$, we measure $\max_{{\mathbf{x}}\in\{{\mathbf{x}}_{1}^{*},...,{\mathbf{x}}_{k}^{*}\}}SSCD({\mathbf{x}}({\mathbf{p}}^{*},{\bm{\epsilon}}_{\theta}),{\mathbf{x}})$. Secondly, we adopt CLIP Score (Hessel et al., 2021) to measure Text-Image alignment between prompt and generated images. Lastly, We adopt an Aesthetic Score (Schuhmann et al., 2022) as the image quality score. While previous works did not measure image quality scores, we observed issues shown in Figure 2. When memorization mitigation methods are applied, we observed that image quality degrades, the rich context generated by the diffusion model is destroyed, or distorted images are formed. An ideal memorization mitigation method should be able to preserve the image generation capabilities of the diffusion model.
@@
To ensure that memorization mitigation methods can be generally applied to diffusion models, we provide two scenarios: the memorized image trigger prompt scenario and the general prompt scenario. First, the memorized image trigger prompt scenario evaluates whether mitigation methods can effectively prevent the generation of memorized images. This scenario uses the memorized image trigger prompts we identified in Section 3. In this scenario, we generate 10 images for each trigger prompt and measure the Top-1 SSCD and the mean values of the Top-3 SSCD. We also measure the proportion of images with SSCD exceeding 0.5. For CLIP Score and Aesthetic Score, we calculate the average value across all generated images. Second, the general prompt scenario ensures that the performance of the diffusion model does not degrade when using prompts other than trigger prompts. We leverage the COCO (Lin et al., 2014) validation set as general prompts. In this scenario, images are generated once per prompt, and the average CLIP Score and Aesthetic Score are measured.
@@
We propose a reference performance for interpreting the performance of mitigation methods. An effective mitigation method should be able to reduce SSCD while maintaining CLIP Score. However, although SSCD is a metric designed to compare the structural similarity of images for copy detection tasks, it inevitably includes semantic meaning due to the self-supervised nature of the trained neural network. On the other hand, the semantic meaning of the trigger prompt should still be reflected in the generated image to maintain CLIP Score even when a mitigation method is applied. Therefore, it is uncertain how much the SSCD between memorized images and generated images can be reduced while maintaining the CLIP Score between trigger prompts and generated images. In this regard, we provide a reference performance to indicate how much SSCD can be reduced while maintaining a high CLIP Score. We assume querying images with trigger prompts via the Google Image API [^3] https://developers.google.com/custom-search as a strong proxy model for the generative model and provide the reference performance based on this approach. Please refer to Appendix for details.
@@ 6 Deeper Analysis into Image Memorization
In this section, we provide a deeper analysis of the memorized images and trigger prompts in MemBench. Firstly, we have found that Stable Diffusion regenerates commercial products currently on sale. While the possibility that diffusion models could memorize commercial images has been suggested (Carlini et al., 2023; Somepalli et al., 2023a), we are the first to confirm this. Unlike the previous studies (Carlini et al., 2023; Webster, 2023) that used image retrieval from LAION to find memorized images, we leverage a Reverse Image Search API to find those, which enable us this verification. As shown in Figure 3.b, Stable Diffusion replicates images of commercially available shirts when given a specific prompt. Figure 3.d further illustrates the replication of layouts; for a commercially sold carpet, all layouts have been reproduced.
Secondly, we explore the cause of image memorization in Stable Diffusion 2, trained on LAION-5B, whose duplicates are removed. Previous works (Somepalli et al., 2023b; Gu et al., 2023) suggested that image memorization issues arise from duplicate images in the training data. Webster et al. (2023) confirmed that the LAION-2B dataset contains many duplicate images likely to be memorized. However, Stable Diffusion 2 still exhibits image memorization issues while reduced. We hypothesize that this memorization arises due to layout duplication. Figure 4 shows the images found by Reverse Image Search API that are memorized by Stable Diffusion. We found that there are often over 100 images on the web with the same layout but different color structures. LAION-5B underwent deduplication based on URLs [^4] https://laion.ai/blog/laion-5b/, but this process may not have removed these images. These layout memorizations are also obviously subject to copyright, posing potential social issues. Additional examples are provided in Appendix.
@@ 7 Evaluation of Image Memorization Mitigation Method on MemBench
Base 0.641 0.605 0.451 0.273 5.25 0.321 5.37
Reference Performance (API search) 0.088 - - 0.310 - - -
RNA(Somepalli et al., 2023b) n = 1 0.479 0.425 0.241 0.270 5.18 0.314 5.34
n = 2 0.389 0.338 0.165 0.270 5.14 0.310 5.33
n = 3 0.329 0.280 0.121 0.267 5.13 0.307 5.30
n = 4 0.287 0.239 0.089 0.264 5.10 0.304 5.29
n = 5 0.254 0.213 0.074 0.262 5.08 0.302 5.28
n = 6 0.228 0.189 0.055 0.258 5.06 0.298 5.24
RTA (Somepalli et al., 2023b) n = 1 0.497 0.446 0.265 0.269 5.20 0.316 5.34
n = 2 0.397 0.347 0.175 0.268 5.19 0.314 5.32
n = 3 0.330 0.285 0.129 0.266 5.17 0.310 5.29
n = 4 0.282 0.240 0.094 0.264 5.15 0.306 5.27
n = 5 0.257 0.217 0.080 0.262 5.14 0.302 5.26
n = 6 0.228 0.190 0.056 0.258 5.10 0.299 5.27
Wen et al. (2024b) l = 7 0.410 0.346 0.134 0.270 5.16 0.321 5.37
l = 6 0.355 0.289 0.089 0.270 5.15 0.321 5.37
l = 5 0.312 0.246 0.059 0.269 5.14 0.321 5.37
l = 4 0.259 0.199 0.035 0.268 5.13 0.321 5.37
l = 3 0.181 0.139 0.015 0.264 5.11 0.321 5.37
l = 2 0.096 0.075 0.001 0.242 4.97 0.321 5.37
Ren et al. (2024) c = 1.0 0.289 0.247 0.083 0.263 5.17 0.316 5.33
c = 1.1 0.283 0.239 0.071 0.260 5.17 0.313 5.31
c = 1.2 0.278 0.232 0.058 0.257 5.15 0.309 5.28
c = 1.3 0.275 0.227 0.050 0.254 5.14 0.304 5.26
Table 3: Performance evaluation of image memorization mitigation methods in MemBench. Please refer to Appendix for the details of hyper-paramters.
In this section, we evaluate image memorization mitigation methods on our MemBench in Stable Diffusion 1. For results of Stable Diffusion 2, please refer to Appendix.
@@
We use Stable Diffusion 1.4 as the base model. The image memorization mitigation methods evaluated include: 1) RTA (Somepalli et al., 2023b), which applies random token insertion to the prompt, 2) RNA (Somepalli et al., 2023b), which inserts a random number between $[0,10^{6}]$ into the prompt, 3) method proposed by Wen et al. (2024b) that applies adversarial attacks to text embeddings, and 4) method proposed by Ren et al. (2024) that rescales cross-attention. Image generation is performed using the DDIM (Song et al., 2021a) Scheduler with a guidance scale of 7.5 and 50 inference steps.
@@ Results.
We present the experimental results in Table 3. As shown in Table 3, for all methods, lowering the SSCD significantly reduces both the CLIP Score and the Aesthetic Score. This indicates a degradation in text-image alignment and image quality. In particular, upon examining images with low Aesthetic Scores, we observe that issues in Figure 2 occur across all methods. While Ren et al. (2024) measured FID, the Aesthetic Score offers a more straightforward way to evaluate individual image quality and better highlight these issues. As reported by Wen et al. (2024b), all methods exhibit a trade-off between SSCD and CLIP Score. Regarding the reference performance obtained via API search, it can be observed that the SSCD can be reduced to 0.088 while maintaining a high CLIP Score. Due to the inherent limitations of the Stable Diffusion baseline model, the CLIP Score cannot exceed 0.273 when mitigation methods are applied. However, mitigation methods should aim to reduce the Top-1 SSCD to around 0.088 while maintaining at least this level of CLIP Score.
To provide a more detailed analysis of each method, we observe that the approach proposed by Wen et al. achieves the best performance in the trade-off between SSCD and CLIP Score. However, to reduce the proportion of images with SSCD exceeding 0.5—indicative of image memorization—to nearly zero, their method still requires a reduction in CLIP Score by 0.025. Given the scale of the CLIP Score, this drop suggests that the generated images may be only marginally related to the given prompts. Moreover, a significant decrease in the Aesthetic Score is also observed. On the other hand, the method proposed by Wen et al. has an additional advantage: it does not result in any performance drop in the general prompt scenario on the COCO dataset, making it the most suitable option for practical applications as of now.
The most recent method proposed by Ren et al. (2024) shows a considerable reduction in the CLIP Score. Even at the lowest hyper-parameter setting ($c=1.0$), the reductions in both CLIP Score and Aesthetic Score are substantial, limiting its general applicability to diffusion models. The most basic approaches, RNA and RTA, show a decrease in CLIP Score by 0.015 at the hyper-parameter setting ($n=6$) that lowers the proportion of images with SSCD exceeding 0.5 to 0.05. This result is expected given the nature of these methods: both attempt to prevent image memorization by adding irrelevant tokens to the prompts. As a result, RNA and RTA are unreliable for application to diffusion models.
@@ Acknowledgments
Ethics statement
Our work introduces a technique for extracting the training data of diffusion models. This could potentially harm the rights of model owners or image copyright holders. Therefore, it is crucial to handle this technique with caution to avoid any infringement issues. For more details, please refer to Appendix.
Reproducibility statement
We provide the code for our training data extraction algorithm, the dataset, and the evaluation in the supplementary material.